# The Language of Approval: Identifying the Drivers of Positive Feedback Online

### v1

@@ 4.1.1. Semantic Style Dimensions (PCs)
We then pick the top-30 and bottom-30 posts by PC scores for each principal component and perform manual inspection as well as cross-validation with GPT-5 model (OpenAI, 2025) (gpt-5-2025-08-07) to interpret what the most aligned and unaligned posts within each dimension indicate. The prompt used can be found in Appendix A. By selecting the posts with the most extreme PC scores along each semantic axis, we hope to capture true signal in terms of what kind of content aligns with these dimensions.
@@ 4.1.4. LIWC Attributes
LIWC Category Odds Ratio $q$-sig Informal language informal residual (other_informal) 0.98 ** Assent 1.02 *** Linguistic Dimensions Conjunctions 0.92 *** Adverbs 1.05 *** Other Grammar Comparisons 1.03 *** Numbers 1.02 * Affective Processes Positive Emotion 1.04 ** Anxiety 1.01 * Sadness 1.01 ** Cognitive Processes cogproc residual (other_cogproc) 1.03 *** Causation 1.02 ** Tentative 0.97 * Perceptual Processes See 0.98 *** Feel 1.01 *** Biological Processes bio residual (other_bio) 0.98 * Ingestion 0.99 ** Drives drives residual (other_drives) 1.03 ** Social Processes Family 1.01 ** Time Orientations Future focus 1.02 *
@@ 4.1.5. Topics:
We found no statistically significant predictors of high score in the set of topic features we modeled. This indicates that no particular topic was unilaterally rewarded across all community using upvotes in our observation window, which might be expected since the communities we study are very diverse in topics.
@@ 4.2. RQ1(b): Do drivers of positive reward differ across feedback outcomes?
[…]
Predictor Odds Ratio $q$-sig Interpretation from exemplars Implication for awards Stylistic Dimensions PC1 1.15 *** Broad engagement prompts over narrow help ↑↑ Engagement framing rewarded PC3 1.05 * Creative or achievement narratives ↑↑ Slight boost for showcases PC4 0.93 *** Dense list-like exposition ↓↓ Dense enumerations penalized PC7 1.05 * Enthusiastic, community-oriented calls ↑ Collective enthusiasm helps PC8 0.97 ** Event-dated or time-stamped content ↓↓ Heavily time-bound posts penalized Surface-level Linguistic Features Question ratio 0.81 *** Question-heavy framing ↓↓ Fewer awards for questions VADER sentiment 1.04 * Overall valence ↑↑ Positive tone slightly rewarded LIWC Attributes Conjunctions 0.97 ** Connectives (e.g., “and,” “whereas”) ↓↓ Over-connected prose penalized Adverbs 0.98 * Degree and manner modifiers ↓↓ Measured descriptiveness fares worse Tentative 0.98 * Hedging (e.g., “maybe,” “perhaps”) ↓↓ Hedging penalized Money 1.03 *** Mentions of money or resources ↑↑ Resource/benefit posts rewarded
@@
Predictor Odds Ratio $q$-sig Implication for gold LIWC Attributes Power 0.94 * ↓↓ Power language disfavored Sadness 1.06 * ↑↑ Empathic responses to sadness Money 1.04 * ↑↑ Resource / benefit posts rewarded
@@ 4.2.1. Agreement and differences with predictors for “score”:
Across channels we see overlaps but also clear divergences. Engagement-oriented framing aligns for both score and awards, with broad, discussion-generating posts being favored in both cases. Similarly, heavily time-bound content and connective-heavy prose is penalized for both outcomes. However, several drivers differ from score. For example the use of adverbs flips its association from score to awards, being positive for score but slightly negative for awards. A slightly more positive overall sentiment matters for awards while it is not a robust causal driver for score. Furthermore, resource-heavy framing shows up in monetarily discretionary channels, with mentions of money being positively associated with both awards and Gold, but are not a key predictor for score. Finally, Gold yields the narrowest set of predictors across all channels, showing a penalty for power and status-driven language and a modest lift for expressions of sadness.
@@ 4.3.1. High score outcome:
[…]
Table with three rows showing interaction effects between newcomer status and linguistic features for high score outcomes. The table has five columns: feature name, interaction odds ratio, statistical significance, veterans odds ratio, and newcomers odds ratio. Readability shows the strongest and most complete effect with interaction odds ratio of 1.18, veterans odds ratio of 1.40, and newcomers odds ratio of 1.65, all with green background coding indicating positive effects. Netspeak and Negation rows show interaction odds ratios of 0.98 and 1.02 respectively, but have dashes in the veterans and newcomers columns indicating non-significant main effects. All three features show single asterisk statistical significance. The readability result demonstrates that while clear writing benefits both groups, newcomers receive disproportionately larger benefits from readable content compared to veteran users. Feature Interaction OR $q$-sig Veterans OR Newcomers OR Readability (Flesch) 1.18 * 1.40 1.65 Netspeak 0.98 * Negation 1.02 *
@@ 4.3.2. Award and gold outcomes:
[…]
Table with three rows showing interaction effects between newcomer status and linguistic features for award outcomes. The table displays five columns showing feature names, interaction odds ratios, statistical significance, veterans odds ratios, and newcomers odds ratios. Future focus shows positive interaction at 1.03 with double asterisk significance, benefiting both veterans at 1.02 and newcomers at 1.05, with green background coding. Anxiety shows negative interaction at 0.97 with single asterisk significance, where veterans benefit at 1.01 but newcomers are penalized at 0.98, creating a mixed color pattern. Informal residual shows negative interaction at 0.97 with single asterisk significance, penalizing both veterans at 0.98 and newcomers more severely at 0.95, both with red background coding. The pattern demonstrates that future-focused language benefits newcomers more than veterans, while anxiety expressions and informal language create differential effects, with newcomers facing greater penalties for informal language and receiving less benefit from anxiety expressions compared to veteran users. Feature Interaction OR $q$-sig Veterans OR Newcomers OR Future focus 1.03 ** 1.02 1.05 Anxiety 0.97 * 1.01 0.98 informal residual 0.97 * 0.98 0.95
@@ 5. RQ2: Predictive Modeling for Proactive Identification of Desirable Comments
We now ask whether the pre-feedback signals studied in RQ1 can support real-time discovery of high quality posts, that is, whether models trained on features observable at posting time can surface promising content that may be underrepresented while votes are still organically accruing. Since the strongest and most stable effects in RQ1 were for the score outcome, we focus on predictive modeling for score only.
@@ 5.2.1. How does our model compare to simple existing desirability measures?
To get a deeper understanding of how well our models perform, we compare the AUC of our model to the prosociality models introduced by Bao et al. (2021), which has been used in prior HCI and social computing work (Goyal et al., 2026; Lambert et al., 2022) as a measure for “desirable” contributions. Moreover, prosociality is considered a key component of desirable contributions on Reddit by moderators (Lambert et al., 2024).
Specifically, we follow the same framework as Goyal et al. (2026) to use the BERT models trained by Bao et al. (2021) to assign each post a supportiveness, agreement, and politeness score, and then normalized all scores. We then used these three scores to synthesized a single-dimensional prosociality measure using Principal Component Analysis (PCA), capturing $\approx 78\%$ of the variance in the original data. We then regressed the binary “high score” label on the post’s synthesized prosociality score.
@@ 5.3. RQ2(b): How does the performance of predictive models vary across communities?
The right panel plot in Figure 4 shows the communities with the top-10 AUC differences (gains and loss) between the local model and the global model evaluated on that community’s test set.
@@
The largest losses (in red) on the other hand occurs in r/PrequelMemes, r/selfie, r/furry_irl, r/PoliticalCompassMemes, r/AnimalsOnReddit, r/WhitePeopleTwitter, r/RedditSessions, r/ShitPostCrusaders, r/moedred, and r/RedditInTheKitchen, with the local models lagging behind the global models by $0.10$ to $0.17$ in terms of AUC. We hypothesize that these subreddits are highly memetic or visual, so text features capture only part of the signal. In such cases the global model benefits from being trained on more data and capturing broader regularities of desirability across Reddit.
@@ 5.3.1. Does distinctiveness play a role?
To explore this pattern further, we test whether out distinctiveness hypothesis is true. To quantify how linguistically “far” a subreddit’s language is from the Reddit‐wide norm, we work with the same 384-dimensional MiniLM embeddings (Reimers and Gurevych, 2019) from our feature set $\mathbf{A}$.
Inspired by Atcheson et al. (2024), for every subreddit $s$ we take a randomly sampled set of $N_{s}=2,000$ (total $N$ across all subreddits) its posts and compute their embeddings $\{\boldsymbol{e}_{s,i}\}_{i=1}^{N_{s}}$. We then compute
@@
Takeaway: Taken together, RQ2 shows that the same pre-feedback linguistic attributes that exhibit causal links to score in RQ1 also carry strong predictive signal. A simple XGBoost detector built on these features achieves $0.654$ AUC in the pooled setting and $0.726$ mean AUC for community specific models, also outperforming existing approaches. We therefore highlight that we can meaningfully rank promising posts at the moment of creation, bridging our causal findings and their practical modeling utility.
@@ 6. RQ3: Alignment with Existing Surveys and Community Guidelines
We now ask whether the cues that causally raise the odds of positive feedback are reflected in what platform users and moderators say they value and in the rules or guidelines that Reddit communities lay out for their users. We split our analysis into two parts, with RQ3(a) comparing our causal estimates with survey-based accounts from users and moderators (Weld et al., 2024; Lambert et al., 2024) and with descriptive inferences from highly upvoted content (Goyal et al., 2026), and RQ3(b) auditing existing subreddit rules and guidelines to test whether they encode these empirically supported cues to enable users to understand these “desirability norms.”
@@ 6.1.1. Comparison with prior surveys (users and moderators):
[…]
Table with 10 rows and 5 columns showing coverage of linguistic attributes in community guidelines across five subreddits. Three symbols indicate coverage levels: explicit checkmark circles, implicit tilde circles, and absent X circles. The first row shows explicit coverage of readability and title clarity across all five subreddits. Rows 2 through 7 covering specificity via numbers, specificity via comparisons, causal framing, future focus, light assent, and concrete social anchors show absent coverage across all subreddits, creating a uniform pattern of X symbols. Row 8 for avoiding toxicity shows mostly explicit coverage with four checkmarks and one tilde symbol. Row 9 for avoiding hedging shows absent coverage across all subreddits. Row 10 for avoiding informal tone shows mixed coverage with two explicit checkmarks and three implicit tildes. The pattern demonstrates that only 3 out of 10 empirically-supported linguistic drivers receive any coverage in community guidelines, with 7 attributes completely absent across all communities, illustrating the systematic gap between causal research findings and actual community guidance practices. Linguistic Attributes r/pcmasterrace r/worldnews r/DnD r/gtaonline r/unpopularopinion Readability / title clarity Specificity via numbers Specificity via comparisons Causal framing (“because,” “effect”) Future focus (“will,” “plan”) Light assent / prosociality Concrete social anchors Avoid toxicity / hostility Avoid heavy hedging / tentative Avoid broad informal tone
@@ 6.1.2. Comparison with descriptive accounts:
Goyal et al. (2026) analyze $16,000$ highly-upvoted comments from $80$ subreddits and, using an LLM pipeline, extract “values” that communities appear to encourage in practice. They report $64$ and $72$ values, extracted from $2016$ and $2022$ data, respectively. Crucially, they find that standard prosociality metrics explain only a small slice of these values, leaving about $82\%$ uncaptured.
Relative to this descriptive picture, our estimates agree that desirability reaches well beyond generic prosociality. The strongest causal levers we identify fit within Goyal et al. (2026)’s “quality” and “accuracy/authenticity” values, and many other effects such as focus future and drives fit within values such as “futurism,” ”progressivism,” “motivation,” and “growth mindset.” Here, our work helps explain why and to what extent those values perform in practice.
Furthermore, we extend the descriptive account by revealing that some intuitive “upvoted” styles are not actually causal drivers of positive feedback. For example, a lot of values extracted by Goyal et al. (2026) did not show up as statistically significant causal predictors or even had slightly negative association in our modeling, e.g., “historical context,” “philosophical,” “dissent.”

### v2

@@ 4.1.1. Semantic Style Dimensions (PCs)
We then pick the top-30 and bottom-30 posts by PC scores for each principal component and perform manual inspection as well as cross-validation with GPT-5 model (OpenAI, 2025) (gpt-5-2025-08-07) to interpret what the most aligned and unaligned posts within each dimension indicate. The prompt used can be found in Appendix A. We found that GPT-5 interpretations matched manual inspection, strengthening our confidence in the dimensions extracted. By selecting the posts with the most extreme PC scores along each semantic axis, we hope to capture true signal in terms of what kind of content aligns with these dimensions.
@@ 4.1.4. LIWC Attributes
LIWC Category Odds Ratio $q$-sig
Informal language
informal residual (other_informal) 0.98 **
Assent 1.02 ***
Linguistic Dimensions
Conjunctions 0.92 ***
Adverbs 1.05 ***
Other Grammar
Comparisons 1.03 ***
Numbers 1.02 *
Affective Processes
Positive Emotion 1.04 **
Anxiety 1.01 *
Sadness 1.01 **
Cognitive Processes
cogproc residual (other_cogproc) 1.03 ***
Causation 1.02 **
Tentative 0.97 *
Perceptual Processes
See 0.98 ***
Feel 1.01 ***
Biological Processes
bio residual (other_bio) 0.98 *
Ingestion 0.99 **
Drives
drives residual (other_drives) 1.03 **
Social Processes
Family 1.01 **
Time Orientations
Future focus 1.02 *
@@ 4.1.5. Topics:
Across varying number of topics ($K$) in our LDA-based topic modeling, we found no statistically significant predictors of high score in the set of topic features we modeled. This indicates that no particular topic was unilaterally rewarded across all community using upvotes in our observation window, which might be expected since the communities we study are very diverse in topics, which may not be adequately captured with 5255-25 topics.
@@
Note: To assess temporal stability of our findings, we split our observation window into two non-overlapping periods (May-July 2020 and July-September 2020) and re-estimated all models independently. The correlation between log-odds ratios across periods is r=0.764(p<0.001)r=0.764~(p<0.001), indicating our findings reflect stable patterns rather than period-specific artifacts (see Appendix B for details).
Predictor Odds Ratio $q$-sig Interpretation from exemplars Implication for awards
Stylistic Dimensions
PC1 1.15 *** Broad engagement prompts over narrow help ↑↑ Engagement framing rewarded
PC3 1.05 * Creative or achievement narratives ↑↑ Slight boost for showcases
PC4 0.93 *** Dense list-like exposition ↓↓ Dense enumerations penalized
PC7 1.05 * Enthusiastic, community-oriented calls ↑ Collective enthusiasm helps
PC8 0.97 ** Event-dated or time-stamped content ↓↓ Heavily time-bound posts penalized
Surface-level Linguistic Features
Question ratio 0.81 *** Question-heavy framing ↓↓ Fewer awards for questions
VADER sentiment 1.04 * Overall valence ↑↑ Positive tone slightly rewarded
LIWC Attributes
Conjunctions 0.97 ** Connectives (e.g., “and,” “whereas”) ↓↓ Over-connected prose penalized
Adverbs 0.98 * Degree and manner modifiers ↓↓ Measured descriptiveness fares worse
Tentative 0.98 * Hedging (e.g., “maybe,” “perhaps”) ↓↓ Hedging penalized
Money 1.03 *** Mentions of money or resources ↑↑ Resource/benefit posts rewarded
@@ 4.2. RQ1(b): Do drivers of positive reward differ across feedback outcomes?
Predictor Odds Ratio $q$-sig Implication for gold
LIWC Attributes
Power 0.94 * ↓↓ Power language disfavored
Sadness 1.06 * ↑↑ Empathic responses to sadness
Money 1.04 * ↑↑ Resource / benefit posts rewarded
@@ 4.2.1. Agreement and differences with predictors for “score”:
Across channels we see overlaps but also clear divergences. Engagement-oriented framing aligns for both score and awards, with broad, discussion-generating posts being favored in both cases. Similarly, heavily time-bound content and connective-heavy prose is penalized for both outcomes. However, several drivers differ from score. For example the use of adverbs flips its association from score to awards, being positive for score but slightly negative for awards. A slightly more positive overall sentiment matters for awards while it is not a robust causal driver for score. Furthermore, resource-heavy framing shows up in monetarily discretionary channels, with mentions of money being positively associated with both awards and Gold, but are not a key predictor for score. Finally, Gold yields the narrowest set of predictors across all channels, showing a penalty for power and status-driven language and a modest lift for expressions of sadness. We view this heterogeneity as a substantive finding and strength of our work. It reveals that free voting mechanisms and paid endorsements encode distinct forms of community value—a distinction obscured by prior surveys that treat all positive feedback as largely equivalent (Lambert et al., 2024; Weld et al., 2024).
@@ 4.3.1. High score outcome:
Table with three rows showing interaction effects between newcomer status and linguistic features for high score outcomes. The table has five columns: feature name, interaction odds ratio, statistical significance, veterans odds ratio, and newcomers odds ratio. Readability shows the strongest and most complete effect with interaction odds ratio of 1.18, veterans odds ratio of 1.40, and newcomers odds ratio of 1.65, all with green background coding indicating positive effects. Netspeak and Negation rows show interaction odds ratios of 0.98 and 1.02 respectively, but have dashes in the veterans and newcomers columns indicating non-significant main effects. All three features show single asterisk statistical significance. The readability result demonstrates that while clear writing benefits both groups, newcomers receive disproportionately larger benefits from readable content compared to veteran users.
Feature Interaction OR $q$-sig Veterans OR Newcomers OR
Readability (Flesch) 1.18 * 1.40 1.65
Netspeak 0.98 *
Negation 1.02 *
@@ 4.3.2. Award and gold outcomes:
Table with three rows showing interaction effects between newcomer status and linguistic features for award outcomes. The table displays five columns showing feature names, interaction odds ratios, statistical significance, veterans odds ratios, and newcomers odds ratios. Future focus shows positive interaction at 1.03 with double asterisk significance, benefiting both veterans at 1.02 and newcomers at 1.05, with green background coding. Anxiety shows negative interaction at 0.97 with single asterisk significance, where veterans benefit at 1.01 but newcomers are penalized at 0.98, creating a mixed color pattern. Informal residual shows negative interaction at 0.97 with single asterisk significance, penalizing both veterans at 0.98 and newcomers more severely at 0.95, both with red background coding. The pattern demonstrates that future-focused language benefits newcomers more than veterans, while anxiety expressions and informal language create differential effects, with newcomers facing greater penalties for informal language and receiving less benefit from anxiety expressions compared to veteran users.
Feature Interaction OR $q$-sig Veterans OR Newcomers OR
Future focus 1.03 ** 1.02 1.05
Anxiety 0.97 * 1.01 0.98
informal residual 0.97 * 0.98 0.95
@@ 5. RQ2: Predictive Modeling for Proactive Identification of Desirable Comments
We now ask whether the pre-feedback signals studied in RQ1 can support real-time discovery of high quality posts, that is, whether models trained on features observable at posting time can surface promising content that may be underrepresented while votes are still organically accruing. Since the strongest and most stable effects in RQ1 were for the score outcome, we focus on predictive modeling only for score to focus on the most important outcome.
@@ 5.2.1. How does our model compare to simple existing desirability measures?
To get a deeper understanding of how well our models perform, we compare the AUC of our model to the prosociality models introduced by Bao et al. (2021), which has been used in prior HCI and social computing work (Goyal et al., 2026a; Lambert et al., 2022) as a measure for “desirable” contributions. Moreover, prosociality is considered a key component of desirable contributions on Reddit by moderators (Lambert et al., 2024).
Specifically, we follow the same framework as Goyal et al. (2026a) to use the BERT models trained by Bao et al. (2021) to assign each post a supportiveness, agreement, and politeness score, and then normalized all scores. We then used these three scores to synthesized a single-dimensional prosociality measure using Principal Component Analysis (PCA), capturing $\approx 78\%$ of the variance in the original data. We then regressed the binary “high score” label on the post’s synthesized prosociality score.
@@ 5.3. RQ2(b): How does the performance of predictive models vary across communities?
The right panel plot in Figure 4 shows the communities with the AUC differences (gains and loss) between the local model and the global model evaluated on that community’s test set. Since our dataset contains 100 subreddits, to extract meaningful insights while keeping interpretations free of bias, we focus on subreddits with both the top-10 gains and top-10 losses as a case study.
@@
The largest losses (in red) on the other hand occurs in r/selfie, r/PrequelMemes, r/ShitPostCrusaders, r/PoliticalCompassMemes, r/moedred, r/AnimalsOnReddit, r/WhitePeopleTwitter, r/RedditInTheKitchen, r/furry_irl, and r/RedditSessions, with the local models lagging behind the global models by $0.10$ to $0.17$ in terms of AUC. We hypothesize that these subreddits are highly memetic or visual, so text features capture only part of the signal. In such cases the global model benefits from being trained on more data and capturing broader regularities of desirability across Reddit.
@@ 5.3.1. Does distinctiveness play a role?
To explore this pattern further and identify ways to mitigate this disparity, we test whether our distinctiveness hypothesis is true. To quantify how linguistically “far” a subreddit’s language is from the Reddit‐wide norm, we work with the same 384-dimensional MiniLM embeddings (Reimers and Gurevych, 2019) from our feature set $\mathbf{A}$.
Operationalizing distinctiveness directly is challenging as it could depend on a variety of factors. To tackle this, we take inspiration from the prior work of Atcheson et al. (2024) to use a distance-based metric for distinctiveness. Specifically, for every subreddit $s$ we take a randomly sampled set of $N_{s}=2,000$ (total $N$ across all subreddits) its posts and compute their embeddings $\{\boldsymbol{e}_{s,i}\}_{i=1}^{N_{s}}$. We then compute
@@
Takeaway: Together, RQ2 shows that the same pre-feedback linguistic attributes that exhibit causal links to score in RQ1 also carry strong predictive signal. A simple XGBoost detector built on these features achieves $0.654$ AUC in the pooled setting and $0.726$ mean AUC for community specific models, also outperforming existing approaches. We therefore highlight that we can meaningfully rank promising posts at the moment of creation, bridging our causal findings and their practical modeling utility.
@@ 6. RQ3: Alignment with Existing Surveys and Community Guidelines
We now ask whether the cues that causally raise the odds of positive feedback are reflected in what platform users and moderators say they value and in the rules or guidelines that Reddit communities lay out for their users. We split our analysis into two parts, with RQ3(a) comparing our causal estimates with survey-based accounts from users and moderators (Weld et al., 2024; Lambert et al., 2024) and with descriptive inferences from highly upvoted content (Goyal et al., 2026a), and RQ3(b) auditing existing subreddit rules and guidelines to test whether they encode these empirically supported cues to enable users to understand these “desirability norms.”
@@ 6.1. RQ3(a): Alignment of Causal Drivers With Prior Surveys and Descriptive Works
A central motivation for this work was to test whether survey-reported community values and patterns observed in highly-upvoted content actually reflect what drives positive feedback on the ground. To examine this alignment rigorously, we employed the quasi-experimental approach in our work. Below, we systematically compare our findings against prior surveys and descriptive work, revealing both where claims align with practice and where they diverge.
@@ 6.1.1. Comparison with prior surveys (users and moderators):
Table with 10 rows and 5 columns showing coverage of linguistic attributes in community guidelines across five subreddits. Three symbols indicate coverage levels: explicit checkmark circles, implicit tilde circles, and absent X circles. The first row shows explicit coverage of readability and title clarity across all five subreddits. Rows 2 through 7 covering specificity via numbers, specificity via comparisons, causal framing, future focus, light assent, and concrete social anchors show absent coverage across all subreddits, creating a uniform pattern of X symbols. Row 8 for avoiding toxicity shows mostly explicit coverage with four checkmarks and one tilde symbol. Row 9 for avoiding hedging shows absent coverage across all subreddits. Row 10 for avoiding informal tone shows mixed coverage with two explicit checkmarks and three implicit tildes. The pattern demonstrates that only 3 out of 10 empirically-supported linguistic drivers receive any coverage in community guidelines, with 7 attributes completely absent across all communities, illustrating the systematic gap between causal research findings and actual community guidance practices.
Linguistic Attributes r/pcmasterrace r/worldnews r/DnD r/gtaonline r/unpopularopinion
Readability / title clarity
Specificity via numbers
Specificity via comparisons
Causal framing (“because,” “effect”)
Future focus (“will,” “plan”)
Light assent / prosociality
Concrete social anchors
Avoid toxicity / hostility
Avoid heavy hedging / tentative
Avoid broad informal tone
@@ 6.1.2. Comparison with descriptive accounts:
Goyal et al. (2026a) analyze $16,000$ highly-upvoted comments from $80$ subreddits and, using an LLM pipeline, extract “values” that communities appear to encourage in practice. They report $64$ and $72$ values, extracted from $2016$ and $2022$ data, respectively. Crucially, they find that standard prosociality metrics explain only a small slice of these values, leaving about $82\%$ uncaptured.
Relative to this descriptive picture, our estimates agree that desirability reaches well beyond generic prosociality. The strongest causal levers we identify fit within Goyal et al. (2026a)’s “quality” and “accuracy/authenticity” values, and many other effects such as focus future and drives fit within values such as “futurism,” ”progressivism,” “motivation,” and “growth mindset.” Here, our work helps explain why and to what extent those values perform in practice.
Furthermore, we extend the descriptive account by revealing that some intuitive “upvoted” styles are not actually causal drivers of positive feedback. For example, a lot of values extracted by Goyal et al. (2026a) did not show up as statistically significant causal predictors or even had slightly negative association in our modeling, e.g., “historical context,” “philosophical,” “dissent.”